A certain metal is 40% tin. How many kilograms of this metal must be mixed with 80 kilograms of a metal that is 70% tin to get a metal that is 50% tin?
Answer by stanbon(75887)

A certain metal is 40% tin. How many kilograms of this metal must be mixed with 80 kilograms of a metal that is 70% tin to get a metal that is 50% tin?
------
Equation:
tin + tin = tin
0.40x + 0.70*80 = 0.50(x+80)
--------
40x + 70*80 = 70x + 50*80
------
-30x = -20*80
x = (2/3)80
---
x = 53 1/3 kg (amt. of 40% metal needed)
=====================
Cheers,
Stan H.
=====================